# Revisiting the Orbital Dynamics of the Hot Jupiter WASP–12b with New Transit Times

### v1

@@ Abstract
In this study, we examine the transit timing variations (TTVs) of the extensively studied hot Jupiter WASP-12b using a comprehensive dataset of 391 transit light curves. The dataset includes 7 new photometric observations obtained with the 1.3 m Devasthal Fast Optical Telescope, the 0.61 m VASISTHA telescope, and the 0.3 m AG Optical IDK telescope, along with 119 light curves from the Transiting Exoplanet Survey Satellite (TESS), 97 from the Exoplanet Transit Database (ETD), 34 from the ExoClock Project, and 134 from previously published sources. To ensure homogeneity and precision, we modeled all 391 light curves and determined their mid-transit times. A detailed transit timing analysis revealed a significant orbital decay rate of $30.31\,\mathrm{ms\,yr^{-1}}$, corresponding to a stellar tidal quality factor of $Q^{\prime}_{\star}=1.61\times 10^{5}$, thereby confirming that the orbit of WASP-12b is indeed decaying rapidly. Furthermore, the computation of model selection metrics ($\chi^{2}_{r}$, BIC, AIC) favors orbital decay as the most likely explanation. However, the presence of an eccentricity above the threshold value allows apsidal precession to remain a viable alternative. We also derived a planetary Love number of $k_{p}=0.66\pm 0.28$, consistent with Jupiter’s value, suggesting a similar internal density distribution. Therefore, while orbital decay is strongly supported, apsidal precession cannot be ruled out as another contributing effect, highlighting the necessity of continued high-precision monitoring to resolve the system’s orbital evolution.
@@ 1 Introduction
The discovery of 51 Pegasi b by Mayor and Queloz in 1995 (M. Mayor & D. Queloz 1995) marked a milestone in exoplanet research, signaling a new era in the study of planets beyond our Solar System. Since this discovery, hot Jupiters, which are gas giant planets with masses similar to Jupiter with short orbital periods (P<10P<10\,~days), have attracted considerable attention and are now among the most extensively characterized exoplanets. For a planet following Keplerian orbit, transits are expected to occur at strictly regular intervals, reflecting a constant orbital period. Deviations from this strict periodicity can indicate the presence of dynamical interactions, such as gravitational perturbations from additional bodies, tidal effects, or relativistic influences (E. Agol et al. 2005; M. J. Holman & N. W. Murray 2005). In particular, tidal interactions in close-in hot Jupiter systems can produce measurable variations in orbital periods (B. Levrard et al. 2009).
@@
Hot Jupiters are especially prone to tidal interactions with their host stars due to their large masses and small orbital separations. Such interactions are now recognized as a key driver of planetary orbital evolution. Through tidal dissipation, i.e., the conversion of tidal forces into internal friction within the stellar interior (P. Hut 1980), these interactions can circularize planetary orbits, enforce spin–orbit synchronization, and, in systems where the planet orbits faster than the stellar rotation rate, trigger tidally driven orbital decay. Previous studies have highlighted the role of stellar age and tidal interactions in shaping the fate of close-in gas-giant planets. For instance, S. Miyazaki & K. Masuda (2023) showed that the occurrence rate of hot Jupiters around Sun-like stars declines with stellar age, implying that close-in gas giants are gradually removed over time, likely due to tidal interactions with their host stars. Similarly, J. H. Hamer & K. C. Schlaufman (2019) argued that tidal interactions can lead to the destruction of hot Jupiters while their host stars are still on the main sequence. These results provide a strong motivation to investigate orbital decay in close-in exoplanetary systems, as it may play a key role in explaining the observed distribution of hot Jupiters. Tidal dissipation is commonly parameterized by the modified stellar tidal quality factor, $Q_{\ast}^{\prime}$. Within this context, a subclass of hot Jupiters with day side temperatures exceeding  2200 K and orbital periods of $\lesssim 2\ \text{days}$, known as ultra-hot Jupiters (UHJs; V. Parmentier et al. 2018), has emerged as particularly favorable systems for exhibiting strong tidal effects.
@@
With observational baselines for many systems now extending beyond a decade, signatures of orbital decay have been reported in an increasing number of hot Jupiters. Candidate systems include HAT-P-19b (S. R. Hagey et al. 2022), HAT-P-32b (S. R. Hagey et al. 2022), HAT-P-51b (L.-C. Yeh et al. 2024), HAT-P-53b (L.-C. Yeh et al. 2024), KELT-9b (J. V. Harre et al. 2023), TrES-1b (S. R. Hagey et al. 2022; E. S. Ivshina & J. N. Winn 2022), TrES-2b (S. R. Hagey et al. 2022; S. Biswas et al. 2024), and TrES-3b, although the evidence remains marginal (S. R. Hagey et al. 2022). For TrES-5b, transit-timing variations consistent with nonlinearity, though not necessarily decay, have been reported (G. Maciejewski et al. 2021; S. R. Hagey et al. 2022; E. S. Ivshina & J. N. Winn 2022; L.-C. Yeh et al. 2024). Compelling evidence of orbital decay has been presented for WASP-4b (L. G. Bouma et al. 2020; S. R. Hagey et al. 2022; J. V. Harre et al. 2023), whereas WASP-32b shows only weak signals (L. Sun et al. 2023). The case of WASP-43b remains debated, with decay reported by L. Sun et al. (2023), but not confirmed by S. R. Hagey et al. (2022). XO-3b has also been suggested as a possible candidate for orbital decay (E. S. Ivshina & J. N. Winn 2022; F. Yang & X. Wei 2022).
@@
To date, the most convincing observational evidence of orbital decay has been obtained for two ultra-hot Jupiters: WASP-12b (K. C. Patra et al. 2017; R. V. Baluev et al. 2019; S. W. Yee et al. 2020) and Kepler-1658b (S. Vissapragada et al. 2022). Their orbital periods are observed to decrease at rates of $29\pm 2$ ms yr-1 and $131\pm 22$ ms yr-1, corresponding to estimated remaining lifetimes of $\sim 3,\mathrm{Myr}$ and $\sim 2.5,\mathrm{Myr}$, respectively. More recently, Ö. Baştürk et al. (2025) reported evidence for a decreasing orbital period in WASP-4b, further strengthening the case for tidal orbital decay among close–in gas giants. Motivated by these findings, we focus our study on WASP-12b, an inflated UHJ discovered by L. Hebb et al. (2009) as part of the Wide Angle Search for Planets (WASP) project. WASP-12b has a mass of $1.47\pm 0.07,M_{\mathrm{J}}$ and a radius of $1.90\pm 0.06,R_{\mathrm{J}}$ (K. A. Collins et al. 2017), and it orbits a late-F type main-sequence star of $\sim 1.4,M_{\odot}$ (K. A. Collins et al. 2017) with a period of $\sim 1.09,\mathrm{days}$. Since its discovery, it has become one of the most intensively studied UHJs, owing to its extreme physical characteristics and its unique role as the first exoplanet with confirmed orbital decay. The primary aim of this work is to present a new transit-timing analysis of WASP-12b by incorporating additional transit observations, thereby extending the temporal baseline available for detecting orbital evolution.
@@
The detection of orbital decay has been enabled by combining long-term transit-timing data from both ground- and space-based surveys. These coordinated efforts have extended the observational timelines of transiting exoplanets, particularly hot Jupiters, over multiple decades, allowing the identification of subtle variations in orbital periods. A major contribution in this area has been made by E. S. Ivshina & J. N. Winn (2022) (hereafter IW22), who compiled a comprehensive catalog of transit timing measurements, including both literature data and observations from the Transiting Exoplanet Survey Satellite. Their database covers 348 systems, of which 240 are hot Jupiters, and the addition of TESS data revealed tentative orbital decay in nine systems not previously identified. Since its launch in 2018, TESS (G. R. Ricker et al. 2014) has played a central role in exoplanetary science, providing high-precision transit timing measurements for nearly all known hot Jupiters orbiting stars brighter than approximately 13th magnitude (D. Huber et al. 2022). This capability has enabled systematic searches for orbital decay across the population of close-in gas giants.
@@
In this study, we re-processed TESS transit timings and incorporated newly observed transits from multiple ground-based observatories worldwide, together with previously published light curves, to extend the temporal baseline of WASP-12b. To further enhance coverage, we also included publicly available data from the Exoplanet Transit Database and the ExoClock Project, which compile contributions from both professional and amateur astronomers. Such combined efforts in regular transit monitoring provide one of the most effective approaches for probing long-term orbital variations in exoplanetary systems. By integrating this extensive dataset with new photometric observations, our work offers a significant improvement over previous studies.
@@
The structure of this paper is as follows: Section 2 describes the target selection. Section 3 presents the observational data collected from various sources, including both space-based and ground-based observations, along with details of data processing and published light curves from the literature. Section 4 outlines the methods and procedures employed to analyze all 391 light curves. Section 5 provides an overview of the transit timing analysis, in which three timing models are fitted to the data. In Section 6, we discuss the results of the TTV analysis. Finally, Section 7 summarizes the conclusions of this work.
2 Target Selection
@@
Previous theoretical work by S. V. Chernov et al. (2017) and N. N. Weinberg et al. (2017) indicated that the observed decay rate in WASP-12b can be explained through gravity wave dissipation, depending on the specific stellar model adopted. However, subsequent analyses by A. J. Barker (2020) were unable to identify any model that satisfies all observational constraints for WASP-12, consistent with the findings of A. Bailey & J. Goodman (2018), who conducted a more extensive parameter survey of the system. More recently, analyses by J. D. Turner et al. (2021) and I. Wong et al. (2022), which included observations from both the primary and extended TESS missions, have provided further support for orbital decay in WASP-12b. Given these conflicting interpretations and the fact that more than a decade has passed since the planet’s discovery, which provides sufficient time to detect orbital decay with a cadence of at least two transits per year, WASP-12b remains an ideal candidate for a detailed TTV study.
3 Observation Log
For WASP-12b, we analyzed photometric transit data from both space and ground based observatories to enhance transit timing precision and extend the temporal coverage. We obtained space based observations from the Transiting Exoplanet Survey Satellite (TESS; G. R. Ricker et al. 2014) and re-processed the data. As for ground based observations, in addition to seven new photometric observations, collected using the 0.61 m VASISTHA Telescope, the 1.3 m Devasthal Fast Optical Telescope (DFOT), and the 0.3 m AG Optical IDK telescope, we have also sourced some ground based light curves from public databases, the Exoplanet Transit Database (ETD; S. Poddaný et al. 2010) and the ExoClock Project (A. Kokori et al. 2023). We obtained 97 high-quality (DQ 1 and 2) light curves from ETD and 34 light curves from the ExoClock Project. Furthermore, 134 additional published light curves were incorporated, with details provided in the following sections. Together, these datasets enable a robust transit timing analysis by combining the high precision of space-based observations with the extended temporal baseline afforded by ground-based monitoring.
3.1 TESS observations
@@
Equipped with four wide-field cameras providing a $24^{\circ}\times 96^{\circ}$ field of view, TESS observes each sector for about 27 days, enabling an almost all-sky survey. During its prime two-year mission (2018–2020), TESS covered $75\%$ of the sky, discovering 66 confirmed exoplanets and identifying over 2100 candidates. These include hot Jupiters, sub-Neptunes, and Earth-sized rocky planets, some within the habitable zone. Notable finds include TOI-700 d (Earth-sized, habitable zone), Pi Mensae c (sub-Neptune), and LHS 3844 b (Earth-sized, key target for JWST studies).
Although its prime mission ended in July 2020, TESS continues in an extended mission, with over 400 confirmed exoplanets and thousands of candidates (as of 2025). Its high-precision transit photometry supports atmospheric studies, orbital dynamics, and JWST target selection, making TESS a cornerstone of modern exoplanet science and a key step toward identifying potentially habitable worlds (G. R. Ricker et al. 2014).
TESS first observed WASP-12 (TIC 86396382) during Sector 20 of its primary mission, covering the period from 2019 December 24 to 2020 January 21 (UT). In this study, we analyzed a total of 391 mid-transit times, including 119 full transit light curves extracted from TESS. The target was observed using Camera 1 with a two-minute cadence across six sectors, 20, 43, 44, 45, 71, and 72, over the time interval from 2019 February 28 to 2023 March 10, with exposure times ranging from 2 to 30 minutes. Observations were performed using both the HLSP Photometer and the TESS photometer mounted on the telescope. Table 1 presents the count of complete transits recorded and the associated data points retrieved from the TESS database for each sector. We have presented the number of full transits observed and the corresponding data points obtained from the TESS database for each sector in Table 1.
The data were downlinked from the spacecraft and processed through the Science Processing Operations Center (SPOC) pipeline at NASA Ames Research Center (J. M. Jenkins et al. 2016). The SPOC pipeline determined the optimal photometric aperture, extracted the light curves, and applied systematic corrections to produce science-ready data products suitable for transit analysis. All final data products are publicly available through the Mikulski Archive for Space Telescopes (MAST).
@@
3.1.1 Data Reduction of high-precision TESS data

### v2

@@ Abstract
In this study, we examine the transit timing deviations of the extensively studied hot Jupiter WASP-12 b using a comprehensive dataset of 391 transit light curves. The dataset includes 7 new photometric observations obtained with the 1.3 m Devasthal Fast Optical Telescope, the 0.61 m VASISTHA telescope, and the 0.3 m AG Optical IDK telescope, along with 119 light curves from the Transiting Exoplanet Survey Satellite (TESS), 97 from the Exoplanet Transit Database (ETD), 34 from the ExoClock Project, and 134 from previously published sources. To ensure homogeneity and precision, we modeled all 391 light curves and determined their mid-transit times. A detailed transit timing analysis revealed a significant orbital decay rate of $31.97\pm 0.80~\mathrm{ms~yr^{-1}}$, corresponding to a stellar tidal quality factor of $Q^{\prime}_{\star}=(1.52\pm 0.038)\times 10^{5}$, thereby confirming that the orbit of WASP-12 b is indeed decaying rapidly. Furthermore, the computation of model selection metrics ($\chi^{2}_{r}$, BIC, AIC) favors orbital decay as the most likely explanation. However, the presence of an eccentricity above the threshold value allows apsidal precession to remain a viable alternative. We also derived a planetary Love number of $k_{p}=0.63\pm 0.089$, consistent with Jupiter’s value, suggesting a similar internal density distribution. In this study, orbital decay is strongly supported, as a plausible cause of the timing deviations observed in WASP-12 system. Continued high-precision monitoring will be essential to further constrain the system’s orbital evolution.
@@ I Introduction
The discovery of 51 Pegasi b by Mayor and Queloz in 1995 (Mayor and Queloz 1995) marked a milestone in exoplanet research, signaling a new era in the study of planets beyond our Solar System. Since this discovery, hot Jupiters, which are gas giant planets with masses similar to Jupiter with short orbital periods (P<10P<10\,~days), have attracted considerable attention and are now among the most extensively characterized exoplanets. For a planet following Keplerian orbit, transits are expected to occur at strictly regular intervals, reflecting a constant orbital period. Deviations from this strict periodicity can indicate the presence of dynamical interactions, such as gravitational perturbations from additional bodies, tidal effects, or relativistic influences (Agol et al. 2005; Holman and Murray 2005). In particular, tidal interactions in close-in hot Jupiter systems can produce measurable variations in orbital periods (Levrard et al. 2009).
@@
Hot Jupiters are especially prone to tidal interactions with their host stars due to their large masses and small orbital separations. Such interactions are now recognized as a key driver of planetary orbital evolution. Through tidal dissipation, i.e., the conversion of tidal forces into internal friction within the stellar interior (Hut 1980), these interactions can circularize planetary orbits, enforce spin–orbit synchronization, and, in systems where the planet orbits faster than the stellar rotation rate, trigger tidally driven orbital decay. Previous studies have highlighted the role of stellar age and tidal interactions in shaping the fate of close-in gas-giant planets. For instance, Miyazaki and Masuda (2023) showed that the occurrence rate of hot Jupiters around Sun-like stars declines with stellar age, implying that close-in gas giants are gradually removed over time, likely due to tidal interactions with their host stars. Similarly, Hamer and Schlaufman (2019) argued that tidal interactions can lead to the destruction of hot Jupiters while their host stars are still on the main sequence. These results provide a strong motivation to investigate orbital decay in close-in exoplanetary systems, as it may play a key role in explaining the observed distribution of hot Jupiters. Tidal dissipation is commonly parameterized by the modified stellar tidal quality factor, $Q_{\ast}^{\prime}$. Within this context, a subclass of hot Jupiters with day side temperatures exceeding  2200 K and orbital periods of $\lesssim 2\ \text{days}$, known as ultra-hot Jupiters (UHJs; Parmentier et al. 2018), has emerged as particularly favorable systems for exhibiting strong tidal effects.
@@
With observational baselines for many systems now extending beyond a decade, signatures of orbital decay have been reported in an increasing number of hot Jupiters. Candidate systems include HAT-P-19 b (Hagey et al. 2022), HAT-P-32 b (Hagey et al. 2022), HAT-P-51 b (Yeh et al. 2024), HAT-P-53 b (Yeh et al. 2024), KELT-9 b (Harre et al. 2023), TrES-1 b (Hagey et al. 2022; Ivshina and Winn 2022), TrES-2 b (Hagey et al. 2022; Biswas et al. 2024), and TrES-3 b, although the evidence remains marginal (Hagey et al. 2022). For TrES-5 b, transit-timing variations consistent with nonlinearity, though not necessarily decay, have been reported (Maciejewski et al. 2021; Hagey et al. 2022; Ivshina and Winn 2022; Yeh et al. 2024; Rothmeier et al. 2025). Compelling evidence for orbital decay has been reported for WASP-4 b (Bouma et al. 2020; Hagey et al. 2022; Harre et al. 2023). More recently, Baştürk et al. (2025) identified a decreasing orbital period in WASP-4 b, further strengthening the case for tidal orbital decay in close-in gas giants; however, Winn and Stefánsson (2025) found that the observed period variations are more consistently explained by light-travel time effects (LTTE) rather than true orbital decay. For WASP-32 b, only weak evidence for orbital decay has been reported (Sun et al. 2023). The status of WASP-43 b remains uncertain, with orbital decay claimed by Sun et al. (2023) but not confirmed by Hagey et al. (2022). XO-3 b has also been proposed as a potential orbital-decay candidate (Ivshina and Winn 2022; Yang and Wei 2022).
@@
To date, the most compelling observational evidence for orbital decay has been obtained for the ultra-hot Jupiter WASP-12 b (Patra et al. 2017; Baluev et al. 2019; Yee et al. 2020). The reality of orbital decay in WASP-12 b has been further confirmed by Winn and Stefánsson (2025). Moreover, high-resolution imaging has revealed that WASP-12 is part of a hierarchical triple stellar system, with two distant M-dwarf companions physically bound to the planet-host star (Bechter et al., 2014). Such higher-order stellar architectures can influence the long-term dynamical and tidal evolution of close-in planets, providing important context for interpreting transit timing variations and orbital decay signals. Motivated by these results, we focus our study on WASP-12 b, an inflated ultra-hot Jupiter discovered by Hebb et al. (2009) as part of the Wide Angle Search for Planets (WASP) project. WASP-12 b has a mass of $1.47\pm 0.07,M_{\mathrm{J}}$ and a radius of $1.90\pm 0.06,R_{\mathrm{J}}$ (Collins et al. 2017), and it orbits a late-F type main-sequence star of $\sim 1.4,M_{\odot}$ (Collins et al. 2017) with a period of $\sim 1.09,\mathrm{days}$. Since its discovery, it has become one of the most intensively studied UHJs, owing to its extreme physical characteristics and its unique role as the first exoplanet with confirmed orbital decay. The primary aim of this work is to present a new transit-timing analysis of WASP-12 b by incorporating additional transit observations, thereby extending the temporal baseline available for detecting orbital evolution.
@@
The detection of orbital decay has been enabled by combining long-term transit-timing data from both ground- and space-based surveys. These coordinated efforts have extended the observational timelines of transiting exoplanets, particularly hot Jupiters, over multiple decades, allowing the identification of subtle variations in orbital periods. A major contribution in this area has been made by Ivshina and Winn (2022) (hereafter IW22), who compiled a comprehensive catalog of transit timing measurements, including both literature data and observations from the Transiting Exoplanet Survey Satellite. Their database covers 348 systems, of which 240 are hot Jupiters, and the addition of TESS data revealed tentative orbital decay in nine systems not previously identified. Since its launch in 2018, TESS (Ricker et al. 2014) has played a central role in exoplanetary science, providing high-precision transit timing measurements for nearly all known hot Jupiters orbiting stars brighter than approximately 13th magnitude (Huber et al. 2022). This capability has enabled systematic searches for orbital decay across the population of close-in gas giants.
@@
In this study, we re-processed TESS transit timings and incorporated newly observed transits from multiple ground-based observatories worldwide, together with previously published light curves, to extend the temporal baseline of WASP-12 b. To further enhance coverage, we also included publicly available data from the Exoplanet Transit Database and the ExoClock Project, which compile contributions from both professional and amateur astronomers. Such combined efforts in regular transit monitoring provide one of the most effective approaches for probing long-term orbital variations in exoplanetary systems.
@@
The structure of this paper is as follows: Section II describes the target selection. Section III presents the observational data collected from various sources, including both space-based and ground-based observations, along with details of data processing and published light curves from the literature. Section IV outlines the methods and procedures employed to analyze all 391 light curves. Section V provides an overview of the transit timing analysis, in which three timing models are fitted to the data. In Section VI, we discuss the results of the transit timing analysis. Finally, Section VII summarizes the conclusions of this work.
II Target Selection
@@
Previous theoretical studies by Chernov et al. (2017) and Weinberg et al. (2017) suggested that the observed orbital decay of WASP-12 b could be driven by dissipation of tidally excited internal gravity waves (IGWs), contingent on the adopted stellar structure. In these models, efficient dissipation arises through wave breaking in a fully radiative stellar core. However, this mechanism requires WASP-12 to be a subgiant star, whereas observational evidence indicates that it is a main-sequence F-type star with a convective core, for which IGW wave breaking is not expected to operate. Subsequent work by Barker (2020), as well as a broader parameter survey by Bailey and Goodman (2018), found no stellar model capable of simultaneously satisfying all observational constraints, highlighting a tension between tidal theory and observations.
More recently, analyses incorporating both primary and extended TESS observations (Turner et al., 2021; Wong et al., 2022) have strengthened the empirical evidence for orbital decay in WASP-12 b. This apparent discrepancy between theory and observations may be alleviated by a newly proposed tidal dissipation mechanism involving the conversion of IGWs into magnetic waves through interaction with a magnetic field generated by a convective core dynamo (see also Duguid et al., 2024). In this framework, efficient tidal dissipation can occur in F-type stars with convective cores over certain evolutionary stages, particularly at older ages. Given these ongoing theoretical developments and the availability of more than a decade of high-precision transit timing data, WASP-12 b remains a prime laboratory for detailed transit timing studies.
III Observation Log
For WASP-12 b, we analyzed photometric transit data from both space and ground based observatories to enhance transit timing precision and extend the temporal coverage. We obtained space based observations from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al. 2014) and re-processed the data. As for ground based observations, in addition to seven new photometric observations, collected using the 0.61 m VASISTHA Telescope, the 1.3 m Devasthal Fast Optical Telescope (DFOT), and the 0.3 m AG Optical IDK telescope, we have also sourced some ground based light curves from public databases, the Exoplanet Transit Database (ETD; Poddaný et al. 2010) and the ExoClock Project (Kokori et al. 2023). We obtained 97 high-quality (DQ 1 and 2) light curves from ETD and 34 light curves from the ExoClock Project. Furthermore, 134 additional published light curves were incorporated, with details provided in the following sections. Together, these datasets enable a robust transit timing analysis by combining the high precision of space-based observations with the extended temporal baseline afforded by ground-based monitoring.
III.1 TESS observations
@@
Equipped with four wide-field cameras providing a $24^{\circ}\times 96^{\circ}$ field of view, TESS observes each sector for about 27 days, enabling an almost all-sky survey. During its prime two-year mission (2018–2020), TESS covered $75\%$ of the sky, discovering 66 confirmed exoplanets and identifying over 2100 candidates. These include hot Jupiters, sub-Neptunes, and Earth-sized rocky planets, some within the habitable zone. Notable finds include TOI-700 d (Earth-sized with habitable zone, Gilbert et al. 2020), Pi Mensae c (super-Earth, Gandolfi et al. 2018), and LHS 3844 b (Earth-sized, a key target for JWST studies, Vanderspek et al. 2019).
Although its prime mission ended in July 2020, TESS continues in an extended mission, with over 400 confirmed exoplanets and thousands of candidates (as of 2025). Its high-precision transit photometry supports atmospheric studies, orbital dynamics, and JWST target selection, making TESS a cornerstone of modern exoplanet science and a key step toward identifying potentially habitable worlds (Ricker et al. 2014).
TESS first observed WASP-12 (TIC 86396382) during Sector 20 of its primary mission, covering the period from 2019 December 24 to 2020 January 21 (UT). In this study, we analyzed a total of 391 mid-transit times, including 119 full transit light curves extracted from TESS. The target was observed using Camera 1 with a two-minute cadence across six sectors, 20, 43, 44, 45, 71, and 72, over the time interval from 2019 February 28 to 2023 March 10, with exposure times ranging from 2 to 30 minutes. Observations were performed using the TESS photometer mounted on the telescope. Table 1 presents the count of complete transits recorded and the associated data points retrieved from the TESS database for each sector. We have presented the number of full transits observed and the corresponding data points obtained from the TESS database for each sector in Table 1.
The data were downlinked from the spacecraft and processed through the Science Processing Operations Center (SPOC) pipeline at NASA Ames Research Center (Jenkins et al. 2016). The SPOC pipeline determined the optimal photometric aperture, extracted the light curves, and applied systematic corrections to produce science-ready data products suitable for transit analysis. All final data products are publicly available through the Mikulski Archive for Space Telescopes (MAST).
@@
III.1.1 Data Reduction of high-precision TESS data